# Distributed Cooperative AI for Large-Scale Eigenvalue Computations Using Neural Networks

### v1

@@ 4.2 Error Analysis
We analyze the error between the estimated eigenvalues and the true eigenvalues of $\mathbf{A}$.
@@
[…]
@@
Under assumptions 1-4, the estimation error satisfies
$\epsilon_{i}^{(k)}\leq\beta\rho^{k}+\gamma,$
where $\beta$ and $\gamma$ are constants dependent on initial conditions and neural network variance.
@@
From the convergence analysis
$\lambda_{i}^{(k)}=\lambda^{*}+\eta^{(k)}+\zeta^{(k)},$
where $\eta^{(k)}$ accounts for consensus error and $\zeta^{(k)}$ accounts for neural network estimation error. We have
$\|\eta^{(k)}\|\leq\rho^{k}e^{(0)},\quad\|\zeta^{(k)}\|\leq\sigma$
Thus
$\epsilon_{i}^{(k)}\leq\rho^{k}e^{(0)}+\sigma=\beta\rho^{k}+\gamma.$
This shows that the error diminishes exponentially with iterations, up to the variance bound of the neural network estimators.∎
Remark 1.
Improving the accuracy of neural network estimators (reducing $\sigma$) directly impacts the final estimation error. Techniques such as increasing training data, enhancing network architecture, and employing regularization can help achieve lower variance.
@@ 4.3 Computational Complexity
We evaluate the computational and communication complexity of the proposed algorithm.
@@
The per-iteration complexity of the distributed cooperative eigenvalue computation algorithm is such that,
  1. (a)
    each agent performs $O(n_{i}^{2})$ operations per iteration, where $n_{i}$ is the size of the submatrix $\mathbf{A}_{i}$.
  2. (b)
@@
Computational Complexity
Notice that each agent $i$ computes the function $f_{\theta_{i}}(\mathbf{A}_{i})$ on its local submatrix $\mathbf{A}_{i}$. Suppose the neural network used by the agent has a fixed architecture (e.g., a fixed number of layers and neurons), then the computational cost of evaluating $f_{\theta_{i}}(\mathbf{A}_{i})$ primarily depends on the size of the input, which is $n_{i}\times n_{i}$. Flattening the matrix $\mathbf{A}_{i}$ to a vector of size $n_{i}^{2}$ results in the computation being proportional to $n_{i}^{2}$ operations. Thus, the computational complexity per iteration for each agent is $O(n_{i}^{2})$.
Communication Complexity
During each iteration, agents communicate with their neighbors by exchanging the scalar values of their current eigenvalue estimates. Since each agent transmits only a single scalar value (the estimated eigenvalue), the communication complexity per iteration is $O(1)$. The minimal communication overhead ensures scalability in the number of agents, as the complexity does not grow with the size of the submatrix or the total number of agents. Thus, the overall per-iteration complexity, combining both computational and communication aspects, is well-balanced and efficient for large-scale distributed computations. ∎The algorithm scales efficiently with the number of agents. Distributing the computation reduces individual computational load, making it suitable for large-scale problems. Communication overhead remains low due to the exchange of scalar values, facilitating implementation over networks with bandwidth constraints.
@@ 4.4 Robustness to Communication Failures
Here, we analyze the algorithm’s robustness in scenarios with intermittent communication failures.
@@
If communication failures are random and the expected communication graph remains connected over time, the distributed cooperative algorithm converges in expectation to the true eigenvalues.
@@
  • the union graph $\mathcal{G}^{\infty}=\bigcup_{t=0}^{\infty}\mathcal{G}_{t}$ is connected
  • there exists a constant $p_{\text{max}}<1$ such that $p_{ij}(t)\leq p_{\text{max}}$ for all $(i,j)$ and $t$
@@
then, the update rule for the eigenvalue estimate $\lambda_{i}^{(k+1)}$ for agent $i$ at time $k+1$ is given by
$\lambda_{i}^{(k+1)}=f_{\theta_{i}}(\mathbf{A}_{i})+\sum_{j\in\mathcal{N}_{i}^{(k)}}w_{ij}^{(k)}(\lambda_{j}^{(k)}-\lambda_{i}^{(k)}),$
where $\mathcal{N}_{i}^{(k)}$ denotes the set of neighbors of agent $i$ in $\mathcal{G}_{k}$, and $w_{ij}^{(k)}$ are the weighting factors satisfying $\sum_{j\in\mathcal{N}_{i}^{(k)}}w_{ij}^{(k)}=1$. The sequence $\{\boldsymbol{\lambda}^{(k)}\}_{k=0}^{\infty}$ where $\boldsymbol{\lambda}^{(k)}=[\lambda_{1}^{(k)},\dots,\lambda_{m}^{(k)}]^{T}$ can be modeled as a stochastic process
@@
Let $\overline{W}^{\infty}=\lim_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\overline{W}_{k}$. Under the assumption that $\mathcal{G}^{\infty}$ is connected, $\overline{W}^{\infty}$ will have a single eigenvalue $1$ with corresponding eigenvector $\mathbf{1}$ and all other eigenvalues strictly less than $1$. The sequence $\{\mathbb{E}[\boldsymbol{\lambda}^{(k)}]\}$ converges to a consensus value $\lambda^{*}$ such that
@@
Implementing redundancy and error-correction mechanisms can further enhance robustness. Adaptive weighting factors $\alpha_{ij}$ based on communication reliability can also improve performance under failures.
@@ 4.5 Comparison with Centralized Methods
We compare the proposed distributed method with traditional centralized eigenvalue computation techniques. For large matrices where $n\gg 1$, the proposed distributed method offers lower computational load per node compared to centralized methods requiring $O(n^{3})$ operations. Moreover, it offers scalability, since adding more agents can handle larger matrices without exponentially increasing computational demands. Finally, the approach also enables faster computations by leveraging concurrent processing. The theoretical analysis confirms that the proposed distributed cooperative AI approach is effective for large-scale eigenvalue computations. It ensures convergence, maintains low error bounds, and offers significant advantages in computational and communication efficiency over traditional methods.
@@ 7 Conclusion
In this paper, we have presented a pioneering distributed neural network-based framework for the estimation of eigenvalues, particularly targeting the smallest eigenvalue of large matrices. Our approach, characterized by its decentralized nature, capitalizes on the synergy between cooperative AI and neural networks to overcome the limitations of conventional eigenvalue computation methods. The proposed system demonstrates not only robustness against communication failures but also scalability, making it highly suitable for large-scale computational problems. Through rigorous theoretical analysis and empirical validation, we have shown that our method consistently achieves high accuracy, with convergence guarantees that surpass existing techniques. The potential applications of this framework extend beyond the specific problem of eigenvalue computation, offering a versatile tool for a wide range of distributed computational tasks. Future research will explore further optimizations and applications in more complex and dynamic environments. This work lays the foundation for a new class of distributed algorithms, heralding significant advancements in computational efficiency and resilience.

### v2

@@ 4.2 Error Analysis
We now analyze the error between the estimated eigenvalues and the true eigenvalues of matrix $\mathbf{A}$.
@@
The estimation error for agent $i$ at iteration $k$ is defined as
$\epsilon_{j}^{(i,k)}=\|\lambda_{j}^{(i,k)}-\lambda_{j}^{*}\|.$
This quantifies the difference between an agent’s estimate and the true eigenvalue.
@@
Under assumptions (i)–(iv) in 4.1 above, the estimation error satisfies
$\epsilon_{j}^{(i,k)}\leq\beta\rho^{k}+\gamma,$
where $\beta$ and $\gamma$ are constants that depend on initial conditions and the variance of the neural network’s estimates.
@@
From the convergence analysis, we have
$\lambda_{j}^{(i,k)}=\lambda_{j}^{*}+\eta_{j}^{(k)}+\zeta_{j}^{(k)},$
where $\eta_{j}^{(k)}$ accounts for the consensus error, and $\zeta_{j}^{(k)}$ accounts for the neural network’s estimation error. The consensus error satisfies $\|\eta_{j}^{(k)}\|\leq\rho^{k}e^{(0)}$, and the neural network estimation error satisfies $\|\zeta_{j}^{(k)}\|\leq\sigma$. Therefore, the total error is
$\epsilon_{j}^{(i,k)}\leq\rho^{k}e^{(0)}+\sigma=\beta\rho^{k}+\gamma.$
This shows that the error diminishes exponentially with iterations, up to the bound imposed by the neural network’s variance.∎
Improving the accuracy of neural network estimators, such as by reducing $\sigma$, directly impacts the final estimation error. Techniques like increasing training data, improving network architecture, and applying regularization can help achieve lower variance and hence better accuracy.
@@ 4.3 Computational Complexity
Finally, we evaluate the computational and communication complexity of the proposed algorithm.
@@
The per-iteration complexity of the distributed cooperative eigenvalue computation algorithm is as follows
  1. 1.
    Each agent performs $O(n_{i}^{2})$ operations per iteration, where $n_{i}$ is the size of the submatrix $\mathbf{A}^{(i)}$.
  2. 2.
@@
  • Computational Complexity: Each agent $i$ computes the function $f_{\theta_{i}}(\mathbf{A}^{(i)})$ on its local submatrix. Given that the neural network has a fixed architecture, the computational cost of evaluating the function primarily depends on the input size $n_{i}\times n_{i}$. The computation is proportional to $n_{i}^{2}$ operations, leading to a per-iteration complexity of $O(n_{i}^{2})$.
  • Communication Complexity: During each iteration, agents communicate their current eigenvalue estimates with their neighbors. Since each agent transmits only a single scalar value (the estimated eigenvalue), the communication complexity per iteration is $O(1)$. This ensures that communication overhead remains minimal and the algorithm scales efficiently with the number of agents.
Thus, the overall per-iteration complexity, combining both computational and communication aspects, is well-balanced and suitable for large-scale distributed computations. Distributing the workload across agents reduces individual computational load, while low communication overhead ensures that the system can operate effectively even in bandwidth-constrained networks.
@@ 4.4 Robustness to Communication Failures
In this section, we analyze the algorithm’s robustness in scenarios where communication between agents is subject to intermittent failures.
@@
If communication failures are random and the expected communication graph remains connected over time, the distributed cooperative eigenvalue computation algorithm converges in expectation to the true eigenvalues.
Definition 3 (Communication Graph).
A communication graph $\mathcal{G}_{t}=(\mathcal{V},\mathcal{E}_{t})$ at time $t$ consists of a set of agents (nodes) $\mathcal{V}$ and a set of communication links (edges) $\mathcal{E}_{t}$ that connect agents. If there is a link between two agents, they can exchange information at time $t$.
Definition 4 (Eigenvalue).
In linear algebra, an eigenvalue is a scalar value that represents how much a matrix stretches or compresses vectors along certain directions (called eigenvectors). In this context, the algorithm computes eigenvalues of matrices distributed across agents.
Definition 5 (Stochastic Process).
A stochastic process is a sequence of random variables evolving over time. In this case, the eigenvalue estimates evolve based on a random communication network.
Definition 6 (Doubly Stochastic Matrix).
A matrix is doubly stochastic if all of its rows and columns sum to 1, meaning the total probability of transitions in a stochastic process is conserved.
Definition 7 (Connected Graph).
A graph is connected if there is a path between any pair of nodes. In the context of communication, it means that agents can eventually exchange information, either directly or indirectly, over time.
Definition 8 (Consensus Value).
A consensus value is a common value that all agents in a distributed system agree on after multiple rounds of communication and updates. In this case, it refers to the converged eigenvalue shared by all agents.
@@
  • The union graph $\mathcal{G}^{\infty}=\bigcup_{t=0}^{\infty}\mathcal{G}_{t}$ is connected.
  • There exists a constant $p_{\text{max}}<1$ such that $p_{ij}(t)\leq p_{\text{max}}$ for all $(i,j)$ and $t$.
@@
The update rule for the eigenvalue estimate $\lambda_{j}^{(i,k+1)}$ for agent $i$ at time $k+1$ is given by
$\lambda_{j}^{(i,k+1)}=f_{\theta_{i}}(\mathbf{A}^{(i)})+\sum_{j\in\mathcal{N}_{i}^{(k)}}w_{ij}^{(k)}(\lambda_{j}^{(k)}-\lambda_{i}^{(k)}),$
where $\mathcal{N}_{i}^{(k)}$ denotes the set of neighbors of agent $i$ in $\mathcal{G}_{k}$, and $w_{ij}^{(k)}$ are the weighting factors satisfying $\sum_{j\in\mathcal{N}_{i}^{(k)}}w_{ij}^{(k)}=1$. The sequence $\{\boldsymbol{\lambda}^{(k)}\}_{k=0}^{\infty}$, where $\boldsymbol{\lambda}^{(k)}=[\lambda_{1}^{(k)},\dots,\lambda_{m}^{(k)}]^{T}$, can be modeled as a stochastic process
@@
Let $\overline{W}^{\infty}=\lim_{K\to\infty}\frac{1}{K}\sum_{k=0}^{K-1}\overline{W}_{k}$. Under the assumption that $\mathcal{G}^{\infty}$ is connected, $\overline{W}^{\infty}$ will have a single eigenvalue equal to 1, with corresponding eigenvector $\mathbf{1}$, and all other eigenvalues strictly less than 1. The sequence $\{\mathbb{E}[\boldsymbol{\lambda}^{(k)}]\}$ converges to a consensus value $\lambda^{*}$ such that
@@
Implementing redundancy and error-correction mechanisms can further enhance robustness. Adaptive weighting factors $w_{ij}$ based on communication reliability can also improve performance under communication failures.
@@ 4.5 Comparison with Centralized Methods
We compare the proposed distributed method with traditional centralized eigenvalue computation techniques. For large matrices where $n\gg 1$, the proposed distributed method offers several advantages over centralized methods
  1. (i)
    Lower Computational Load: In centralized methods, the computational cost for eigenvalue computation scales as $O(n^{3})$. In contrast, in the distributed method, each agent handles a smaller submatrix $\mathbf{A}^{(i)}$, reducing the computational load per agent to $O(n_{i}^{2})$.
  2. (ii)
    Scalability: The distributed method is scalable, as adding more agents increases the overall system capacity without exponentially increasing the computational demands on any single agent.
  3. (iii)
    Concurrency: The distributed nature of the algorithm allows concurrent processing across multiple agents, speeding up the computation in large-scale problems.
The theoretical analysis confirms that the proposed distributed cooperative method is effective for large-scale eigenvalue computations. It ensures convergence, maintains low error bounds, and offers significant advantages in computational and communication efficiency over traditional methods.
@@ 7 Conclusion
In this paper, we have introduced a novel distributed neural network-based framework for estimating eigenvalues, with a particular focus on the smallest eigenvalue of large matrices. The decentralized nature of our approach leverages cooperative AI and neural networks to address the limitations of traditional eigenvalue computation methods. The communication graph, along with the convergence analysis, demonstrates the robustness and efficiency of the proposed system in a distributed environment. Our empirical results, particularly the convergence plot and the close alignment between the true and estimated smallest eigenvalues, confirm the accuracy and precision of the method. Despite minor deviations typical of distributed systems, the agents’ estimates converge to values that closely approximate the true eigenvalue, underscoring the algorithm’s reliability. Moreover, the system is resilient to communication failures and scalable for large-scale computational tasks, as evidenced by the performance over 100 iterations and the network’s high degree of connectivity. Theoretical and practical findings validate the proposed framework’s effectiveness, offering a promising tool for solving large-scale eigenvalue problems in distributed settings. Future work will focus on optimizing the algorithm further and expanding its application to more complex computational environments. This research opens new avenues for distributed algorithms, advancing computational efficiency and robustness in large-scale systems.